### v2

# Online SLA Decomposition: Enabling Real-Time Adaptation to Evolving Systems

Abstract
When a network slice spans multiple domains, each domain must uphold the End-to-End (E2E) Service Level Agreement (SLA) associated with the slice. This requires decomposing the E2E SLA into partial SLAs for each domain. In a two-level network slicing management system with an E2E orchestrator and local controllers, we propose an online learning-decomposition framework that dynamically updates risk models using recent feedback. This approach utilizes online gradient descent and FIFO memory buffers to enhance stability and robustness. Our empirical study shows the proposed framework outperforms state-of-the-art static methods, offering more accurate and resilient SLA decomposition under varying conditions and sparse data.
Index Terms:
network slicing, service level agreement, risk model, quality of service, deep neural network, online learning
I Introduction
The fifth generation (5G) of mobile communication technology introduced a versatile, multi-service network designed to support a wide range of vertical industries with a diverse set of service requirements. In 5G and beyond, network slicing plays a pivotal role by enabling the establishment and management of multiple End-to-End (E2E) logical networks. These slices are built on shared infrastructure and are specifically customized to meet the particular requirements of a given service, which is outlined in Service Level Agreements (SLAs). SLAs function as contracts between service providers and tenants, defining the expected Quality of Service (QoS) through well-defined, measurable benchmarks known as Service-Level Objectives (SLOs). These objectives encompass various performance metrics such as data throughput, latency, reliability, and security, among others. A single network slice may traverse multiple segments of the network, including access, core, and transport networks, and it may involve collaboration between different operators and infrastructure providers. To ensure that the service meets the agreed-upon SLOs across these domains, it is essential to adjust the service parameters accordingly. As a result, the E2E SLA linked to a network slice must be broken down into specific SLOs for each domain, allowing for effective resource allocation within each segment. The authors in [1] highlight the importance of E2E SLA decomposition for resource allocation, while [2] indicates that AI-assisted SLA decomposition is key to automating complex 6G business. Following the scenarios in [3, 4], in this paper we consider a two-level management architecture with an E2E service orchestrator handling network service lifecycle management, and local domain controllers managing slice instantiation within their domains (Fig. 1). The orchestrator determines the SLA decomposition for incoming service requests, while domain controllers handle admission control and resource allocation. We assume the orchestrator lacks real-time knowledge of the infrastructure state but has access to historical feedback (i.e., request acceptance or rejection) from each domain. This allows the orchestrator to make informed decisions using domain-specific risk models based on the available data.
Figure 1: Network slicing management and orchestration system, and the SLA decomposition module.
Several SLA decomposition methods employing heuristics have been studied [1]. Authors in [5] present an E2E SLA decomposition system that applies supervised machine learning to break down E2E SLAs into access, transport, and core SLOs. In our previous work [3, 4], we tackle the problem with a two-step approach, which is a combination of machine learning and optimization-based solution.
While the approaches in [3, 4] have demonstrated success, they do not consider the inherent dynamicity of the system, which is critical for precise SLA management across network domains. This dynamicity is shaped by several factors, including variations in traffic intensity, shifts in user behavior, and fluctuating network conditions. As these factors evolve, the dynamics within each domain also change, potentially affecting the decision-making process of domain controllers. To address this gap, we propose an online learning-decomposition framework on top of [4], specifically tailored for SLA management in dynamic, multi-domain environments.
The contributions of this paper are summarized as:
  1. 1.
    We propose an online learning-decomposition framework that continuously updates risk models based on real-time feedback from domain controllers, enabling adaptation to dynamic conditions.
  2. 2.
    We leverage First In First Out (FIFO) memory buffers for data management to improved robustness and performance.
  3. 3.
    We conduct an ablation study to examine the contribution and importance of individual component of the proposed framework, and compare with the state-of-the-art offline learning approach.
The rest of this paper is organized as follows; Section II presents background information and problem formulation on the SLA decomposition. In Section 2 we present the proposed online learning-decomposition framework. Section IV details the experimental setup, followed by Section V, where we assess the performance of the proposed framework. Finally, Section VI presents our conclusions.
II Problem Description
II-A SLA Decomposition and Risk Models
An E2E SLA, denoted by $s_{e2e}$, is a collection of Service Level Objectives (SLOs) linked to specific performance indicators. The SLO vector outlines the SLA’s performance requirements in a sequential manner. For example, an SLA encompassing E2E delay and throughput is expressed as $s_{e2e}=(\tau_{e2e},\theta_{e2e})$. This implies that the network slice must operate in a way that meets constraints imposed by $\tau_{e2e}$ for delay and $\theta_{e2e}$ for throughput, ensuring $\tau\leq\tau_{e2e}$, and $\theta\geq\theta_{e2e}$. Considering a network slice distributed across $N$ domains (where n ranges from 1 to $N$), we introduce $s_{n}$ to represent the SLOs of the $n$-th domain. The relationship between individual domain SLOs and the overall E2E objective $s_{e2e}$ is defined by $s_{e2e}=G(s_{1},s_{2},...,s_{N})$. For instance, the E2E delay is the sum of all delays for the involved domains, while the E2E throughput is determined by the lowest throughput across all domains. Mathematically, this is represented as $\tau_{e2e}=\Sigma_{n=1}^{N}\tau_{n}$ and $\theta_{e2e}=\min\{\theta_{1},\theta_{2},...,\theta_{N}\}$.
We can model the ability of a domain to support a partial SLA $s_{n}$ with a risk model, and the risk models of all involved domains can then be used in the SLA decomposition process. The risk model is defined as $-\log P_{n}(s_{n})$, where $P_{n}(s_{n})$ represents the probability that a request in the $n$-th domain with SLOs $s_{n}$ is accepted. Under the assumption of independent decision-making by each domain, the overall E2E acceptance probability is calculated as the product of the individual acceptance probabilities of all involved domains. Therefore, the E2E decomposition can be formulated as an optimization problem that minimizes the overall risk in objective (1) under the constraints (2):
$-\sum_{n}^{N}\log P_{n}(s_{n})$ (1)
$s_{e2e}=G(s_{1},s_{2},...,s_{N})$ (2)
II-B Determining Neural Network-based Risk Models
We determine the risk model per domain with a parameterized Neural Network (NN) $\mathcal{F}$ [4], where the probability $P_{n}(s_{n})$ is modelled as $\mathcal{F}_{n}(s_{n})$. When a domain is presented with a new service request with specific SLOs, denoted by $s$ (domain subscript omitted for simplicity), a controller must determine whether to accept or reject the request. This decision depends not only on the requested SLOs but also on the current infrastructure state, represented by $\omega$. The infrastructure state encompasses factors such as link and server utilization, network hop delays, and available backup paths, etc. Therefore, the decision to accept or reject a service request depends jointly on the SLOs $s$ and infrastructure state $\omega$. While the domain controller has granular visibility of the infrastructure state, the orchestrator lacks this level of detail. Accordingly, although the controller’s decision-making process is deterministic, the orchestrator perceives it as stochastic due to the unknown infrastructure state $\omega$.
Nevertheless, the acceptance probability $P$ can be estimated by analyzing the domain controller’s responses to past requests. Given a set of $K$ responses to previous requests $\{(x_{1},y_{1}),(x_{2},y_{2}),...,(x_{K},y_{K})\}$, each represented by a proposed SLO $x$ and its corresponding acceptance decision $y$ (0 for rejection, 1 for acceptance), we can model the acceptance probabilities for SLO vectors with parameterized neural networks $\mathcal{F}$ by maximizing the overall likelihood on the dataset:
$\sum_{i=1}^{K}[y_{i}*\log(\mathcal{F}(x_{i}))+(1-y_{i})*\log(1-\mathcal{F}(x_{i}))].$ (3)
Furthermore, the acceptance probability exhibits a partial ordering relation [3], which incorporates the concept of SLA strictness, i.e., given a set of $K$ SLOs $S=\{x_{1},x_{2},...,x_{K}\}$, the acceptance probability has the following property:
$\forall x_{i},x_{j}\in S,P(x_{i})\leq P(x_{j})\quad\textrm{if}\quad x_{i}\preceq x_{j},$ (4)
The property indicates that a stricter SLO $x_{i}$ is less likely to be accepted compared to $x_{j}$. The authors in [4] proposed several effective methods to bake this property into NN-based risk models without incurring any architectural constraints.
II-C Dynamicity
The dynamicity of the system is crucial for accurately managing SLAs in network domains. Multiple factors contribute to this dynamicity, including traffic intensity fluctuations, changes in user behavior, varying network conditions and security threats. As these factors change, the acceptance probability of requests within a domain varies, impacting overall network performance. Particularly, the acceptance probability becomes time-dependent. At a discrete time step $t$, the acceptance probability (in a single domain) of the SLA request $s$ is denoted as $P_{t}(s)$, which is influenced by the state $\omega_{t}$ of the domain, as stated in Section II-B. Given that $\omega_{t}$ evolves over time, the acceptance probability $P_{t}(s)$ also varies, even for the same SLOs. This behavior necessitates continuous learning and adaptation of the risk models. Dynamic systems require periodic updates to risk models based on the recent feedback to ensure these models remain relevant and accurate. In the next section, we introduce a novel framework that allows the system to adapt to varying conditions and maintain near-optimal performance.
III Approaches
Figure 2: Illustration of RADE framework.
In this section, we present the proposed framework Real-time Adaptative DEcomposition (RADE).
III-A Two-step Decomposition with NN-based Risk Models
A two-step decomposition approach was proposed in [3, 4]. A dedicated risk model is learned first for each domain, given the historical feedback (i.e., proposed SLAs and their corresponding acceptance/rejection decisions made by controllers) from domain controllers. Once the risk models are built, an optimization proceeds to search for the decomposition that maximizes the E2E acceptance probability, as formulated in (1) and (2). However, in the previous work, the risk models are trained once and the weights are kept fixed for all future SLA decompositions. This approach is impractical for real-world systems, which are typically dynamic and time-dependent. The static models often lead to sub-optimal performance. We describe our extended solution to tackle this problem in the next subsection.
III-B Online Learning Framework
To capture the dynamic nature of the environment, where the decision-making process of controllers evolves over time, it is essential to constantly update risk models based on the recent feedback. To this end, we propose an online learning-decomposition framework RADE, which is capable of running stable update as well as providing resilience against noisy samples. The following are the key components of RADE: Base model. Following the design in [4], the base model is an NN. To account for monotonicity, we specifically choose Absolute Weight Transformation (AWET) model that shows prominent performance while requiring the least training effort, as suggested in [4].
Online update. Unlike traditional static models, which are trained once and applied indefinitely, our approach involves periodic updates to the model based on the most recent feedback collected within each discrete time step. As illustrated in Fig. 2, the loop begins with a base model and employs simple Online Gradient Descent (OGD) [6] to perform updates. The continuously updated model is then used for real-time decomposition, ensuring that the system adapts promptly to the latest conditions.
FIFO memory buffer. Updating the model solely based on the most recent observations can lead to instability. For instance, the model may overfit when the feedback data is sparse, or learning may be compromised if feedback data contains errors. To mitigate these issues, we propose using a FIFO buffer with finite capacity for storing feedback. The FIFO buffer ensures a more stable and reliable learning process by maintaining a portion of historical feedback alongside all recent feedback [7]. The limited capacity is necessary to ensure that outdated information is discarded, allowing the dataset to remain current and relevant for ongoing learning. This component helps prevent overfitting by providing a more diverse set of training samples and safeguards against the detrimental effects of occasional corrupted feedback.
Online decomposition. Besides the learning loop, the inference loop is also running in parallel. Upon receiving a new request, the associated E2E SLA will be decomposed with the latest risk models. The optimization-based decomposition follows the one proposed in [4]. By leveraging up-to-date risk models, the inference loop ensures that the decomposition accurately reflects the current state of the corresponding domains.
Input: Step size $\eta$, risk models $\mathcal{F}_{\theta}$, Memory buffer $\mathcal{M}$
for time step $t=1,2,\ldots,T$ do
      In parallel:
       begin Learning:
             Receive feedback $\mathcal{K}_{t}=\{k_{1},k_{2},...,k_{m}\}$
             $\mathcal{M}\leftarrow\text{FIFO-PUSH}(\mathcal{M},\mathcal{K}_{t})$
             $\theta\leftarrow\theta-\eta\nabla\mathcal{F}_{\theta}(\mathbf{\mathcal{M}})$ // multiple runs
       end
      begin Inference:
             Receive requests $\mathcal{R}_{t}=\{r_{1},r_{2},...,r_{n}\}$
             with associated SLAs $\mathcal{S}_{e2e}$
             $\mathcal{S}_{partial}\leftarrow decompose(\mathcal{F}_{\theta},\mathcal{R}_{t},\mathcal{S}_{e2e})$
             return $\mathcal{S}_{partial}$
       end
end for
Algorithm 1 RADE framework
By updating the risk models online, we ensure that the decomposition remains responsive and resilient to changes. This continuous adaptation process not only enhances the accuracy of the model but also improves the overall performance and reliability of the system. Alg. 1 details the steps of the online update and decomposition mechanism. In the subsequent sections, we will present the simulation environment for evaluations, followed by empirical results demonstrating its effectiveness.
IV Experimental Setup
IV-A Simulation Environment
We follow the analytic model and data generation process proposed in [3, 4] to generate data for three domains. This analytic model maps a decomposition assignment to a probability, indicating how possible the given assignment will be accepted by the current domain controllers. To introduce dynamicity into domain controller’s decision-making process over time, we assume that the acceptance probability is inversely proportional to the current traffic intensity. Specifically, we define a time-dependent factor:
$\lambda_{t}=\frac{1}{2}\left(\frac{\sin(2\pi*t)}{N}+1\right)*0.9+0.1,$ (5)
which represents the traffic intensity at time step $t$ over total number of steps $N$, as used in [8]. We then upgrade the form factor $\alpha$ proposed in [3], which models the probability distribution of the current load on the system, to $\alpha^{\prime}=\frac{\alpha}{\lambda_{t}}$, thereby accounting for the impact of varying load on the acceptance probability. This adjustment reflects the changing traffic conditions over time. Note that the form factor $\alpha$ controls the level of acceptance probability: a larger $\alpha$ results in a higher acceptance probability for the same SLA request.
IV-B Evaluation scenarios and metrics
Average acceptance probability over time. We run Alg. 1 within the simulation environment detailed in Section IV-A. Given the response to each decomposition assignment is represented as a probability value $\in[0,1]$, the E2E acceptance probability at each time step $t$ is calculated as the product of the individual acceptance probabilities across all three domains. The average acceptance probability is then reported over the entire simulation period:
$p_{avg}=\frac{1}{T}\sum_{t=1}^{T}\frac{1}{M_{t}}\sum_{m=1}^{M_{t}}\prod_{d=1}^{D}P_{d,t}^{am}(s_{m,d,t}),$ (6)
where $T$ denotes the number of total time steps, $M_{t}$ is the number of requests at time $t$, $D$ is the number of involved domains (which is set to 3 in this paper), $P^{am}$ represents the analytic model, and $s_{m,d,t}$ is the decomposed partial SLA of the $m$-th request that is assigned to the $d$-th domain at time $t$. The number of requests $M_{t}$ at time $t$ is sampled from the Poisson distribution with $\lambda=\lambda_{t}$ described in (5). The E2E SLA for each request is given as $(\tau_{e2e},\theta_{e2e})$, where $\tau_{e2e}$ and $\theta_{e2e}$ are sampled uniformly from [90ms, 110ms] and [0.4Gbps, 0.6Gbps], respectively.
Resilience test. To assess the framework’s resilience against corrupted feedback labels, we perform a resilience test on top of the aforementioned test, where each feedback has a corruption probability $p_{c}$ of being corrupted (i.e., the request is always rejected), which can result from issues like network delays, transient errors, or misconfiguration in domain controllers. Specifically, we track how the model’s accuracy is affected as $p_{c}$ increases. This resilience test offers crucial insights into the effectiveness of the use of FIFO memory buffer mentioned in Section 2-B, and the overall robustness of our framework under challenging conditions.
IV-C Configurations
To evaluate the contribution of each component proposed in Section 2-B, we perform an ablation study by incrementally adding improvements at three stages, with each stage forming a distinct comparison method. Furthermore, two additional methods (Random and OPT) are included as benchmarks.
Random. The Random method does not employ any risk models and instead decomposes each incoming request’s E2E SLA uniformly at random. This method is used as a baseline to verify the effectiveness of the proposed methods.
Static. The Static method involves a one-time training of risk models using feedback collected from a single prior run with the Random method. The weights of these risk models are then fixed and applied to all subsequent decompositions over time. The configuration of the NN-based risk models is consistent with that described in [4]: a 3-layer multilayer perceptron (MLP), with 8 neurons each. The hyperbolic tangent (Tanh) activation function and Batch Normalization (BN) are applied for hidden layers in the order of linear-Tanh-BN.
RADE/RADE*. RADE represents the full method described in Section 2-B, whereas RADE* is a variant of RADE that omits the FIFO memory buffer. In RADE*, risk models are updated using only the most recent feedback. Both RADE and RADE* utilize the Static method to initialize the risk models.
OPT. An exhaustive search is conducted at every time step to find the decomposition assignment that yields the largest E2E acceptance probability. This method provides the maximum theoretical performance achievable.
V Results
Fig. 3 presents the average E2E acceptance probability across different arrival rates for four methods: Static, RADE*, RADE, and OPT. The arrival rates vary between 0.3, 0.5, and 0.7, representing different traffic intensities in the system. Across all arrival rates, the Static method consistently exhibits the lowest average E2E acceptance probability. This suggests that the Static method’s one-off trained risk models are less adaptable to varying traffic conditions, leading to sub-optimal performance, particularly under higher traffic loads. RADE* shows improved performance over the Static method, which indicates that dynamically updating the risk models, even without the FIFO memory buffer, leads to a better adaptability than the Static approach. The RADE method, which includes the FIFO memory buffer for maintaining historical feedback, outperforms both Static and RADE*, highlighting the importance of the use of the FIFO buffer in enhancing the robustness and stability of the framework, particularly under varying traffic conditions. It is interesting to see that the performance of RADE* and RADE becomes nearly identical at higher arrival rates because, under heavy traffic, the system receives a large volume of recent feedback. This plenty of fresh data diminishes the impact of the FIFO memory buffer in RADE, as the most recent observations dominate the learning process. As expected, the OPT method achieves the highest performance, which serves as a theoretical upper bound. In contrast, the Random method performs significantly worse than the other approaches in general, with results of 0.42, 0.38, and 0.35 for arrival rates of 0.3, 0.5, and 0.7, respectively. Due to the consistently low performance of the Random method, we exclude its results from all subsequent figures to focus on the more meaningful comparisons.
Fig. 4 shows the E2E acceptance probability over time of one run for three methods along with the corresponding arrival rate (the dashed line in red on the secondary y-axis). The Static method exhibits significant fluctuations and lower performance in general. RADE* performs well initially, but its performance degrades sharply when the arrival rate is low (around the $200$-th time step). This drop is likely due to its reliance on only the most recent feedback, making it unstable during periods of low traffic that only sparse feedback are available. RADE consistently outperforms both Static and RADE*, maintaining higher and more stable acceptance probabilities across different arrival rates. The stability and adaptability of RADE are particularly evident during the low arrival rate period, where RADE* struggles.
Fig. 5 illustrates the average E2E acceptance probability for the RADE* and RADE methods under varying corruption rates (0.1, 0.2, and 0.3). As the rate increases, the performance of RADE* significantly deteriorates, while RADE consistently maintains a higher acceptance probability across all corruption rates. This figure clearly demonstrates again that RADE is more resilient to corrupted feedback compared to RADE*, due to the use of the FIFO memory buffer.
Figure 3: Average E2E acceptance probability vs. arrival rate.
Figure 4: One run of E2E acceptance probability over time.
Figure 5: Average E2E acceptance probability vs. corruption rate at arrival rate=0.5.
VI conclusion
This paper has introduced an online learning-decomposition framework for SLA management in dynamic multi-domain environments. The proposed framework updates domain-specific risk models in real-time based on recent feedback, hence effectively addressing the challenges posed by the dynamicity of real-world systems. With the use of online update and FIFO memory buffers, the proposed framework enhances the stability and robustness of SLA decomposition, ensuring that the system is able to adapt promptly to changes in network conditions. Empirical results demonstrate that our framework outperforms state-of-the-art static methods, offering a more resilient and accurate decomposition of SLAs, even with limited data availability. In future work, we will investigate the potential of Deep Reinforcement Learning (DRL) to further enhance the SLA decomposition process. DRL could be a promising choice for this task due to its ability to learn optimal long-term policies through interaction with dynamic environments, which makes it an attractive option for improving SLA management in next-generation networks.

### v3

@@
 Abstract
-When a network slice spans multiple domains, each domain must uphold the End-to-End (E2E) Service Level Agreement (SLA) associated with the slice.
-This requires decomposing the E2E SLA into partial SLAs for each domain.
-In a two-level network slicing management system with an E2E orchestrator and local controllers, we propose an online learning-decomposition framework that dynamically updates risk models using recent feedback.
-This approach utilizes online gradient descent and FIFO memory buffers to enhance stability and robustness.
-Our empirical study shows the proposed framework outperforms state-of-the-art static methods, offering more accurate and resilient SLA decomposition under varying conditions and sparse data.
+When a network slice spans multiple technology domains, it is crucial for each domain to uphold the End-to-End (E2E) Service Level Agreement (SLA) associated with the slice.
+Consequently, the E2E SLA must be properly decomposed into partial SLAs that are assigned to each domain involved.
+In a network slice management system with a two-level architecture, comprising an E2E service orchestrator and local domain controllers, we consider that the orchestrator has access solely to historical data regarding the responses of local controllers to previous requests, and this information is used to construct a risk model for each domain.
+In this study, we extend our previous work by investigating the dynamic nature of real-world systems and introducing an online learning-decomposition framework to tackle the dynamicity.
+We propose a framework that periodically updates the risk models based on the most recent feedback.
+This approach leverages key components such as online gradient descent and FIFO memory buffers, which enhance the stability and robustness of the overall process.
+Our empirical study on an analytic model-based simulator demonstrates that the proposed framework outperforms the state-of-the-art static approach, providing more accurate and resilient SLA decomposition even under varying conditions and limited data scenarios.
 Index Terms:
@@
 These objectives encompass various performance metrics such as data throughput, latency, reliability, and security, among others.
-A single network slice may traverse multiple segments of the network, including access, core, and transport networks, and it may involve collaboration between different operators and infrastructure providers.
+A single network slice may traverse multiple segments of the network, including (radio) access, transport, and core networks, and it may involve collaboration between different operators and infrastructure providers.
 To ensure that the service meets the agreed-upon SLOs across these domains, it is essential to adjust the service parameters accordingly.
 As a result, the E2E SLA linked to a network slice must be broken down into specific SLOs for each domain, allowing for effective resource allocation within each segment.
-The authors in [1] highlight the importance of E2E SLA decomposition for resource allocation, while [2] indicates that AI-assisted SLA decomposition is key to automating complex 6G business.
-Following the scenarios in [3, 4], in this paper we consider a two-level management architecture with an E2E service orchestrator handling network service lifecycle management, and local domain controllers managing slice instantiation within their domains (Fig.
-1).
+The problem has been described in [1, 2, 3].
+In [1], authors suggested decomposing the end-to-end requirement to the transport network requirement is one of the key issues in network slice requirement mapping.
+Authors in [2] outlined how slices are managed through lifecycle automation, orchestration, and real-time monitoring to ensure SLA compliance across different domains.
+In [3], authors elaborate that E2E QoS depends on SLA parameters and that transport resources must align with the heterogeneous QoS needs of different slices.
+The authors in [4] highlighted the importance of E2E SLA decomposition for resource allocation, while [5] indicates that AI-assisted SLA decomposition is key to automating complex 6G business.
+Following the scenarios in [6, 7], in this paper we consider a two-level management architecture with an E2E service orchestrator handling network service lifecycle management, and local domain controllers managing slice instantiation within their domains, as shown in Fig.
+1.
 The orchestrator determines the SLA decomposition for incoming service requests, while domain controllers handle admission control and resource allocation.
@@
 Figure 1: Network slicing management and orchestration system, and the SLA decomposition module.
-Several SLA decomposition methods employing heuristics have been studied [1].
+Some studies [8, 9, 10] have proposed prediction-based approaches for SLAs management, although they do not explicitly address E2E SLAs decomposition problem.
+In [8], authors proposed the use of a mapping layer, which supervises the network over a service area and manages the allocation of radio resources to slices to guarantee their target service requirements.
+Authors in [9] proposed a SLA-constrained optimization using Deep Learning (DL) to estimate the required resources based on the traffic per slice.
+Another work [10] employs a context approach using graph representations for SLA violations prediction in Cloud Computing.
+Moreover, several SLA decomposition methods employing heuristics have been studied [4].
 Authors in [11] present an E2E SLA decomposition system that applies supervised machine learning to break down E2E SLAs into access, transport, and core SLOs.
@@
   3. 3.
-    We conduct an ablation study to examine the contribution and importance of individual component of the proposed framework, and compare with the state-of-the-art offline learning approach.
+    We conduct an ablation study to examine the contribution and importance of individual component of the proposed framework, and compare with the state-of-the-art offline learning approach, followed by a complexity analysis.
 The rest of this paper is organized as follows; Section II presents background information and problem formulation on the SLA decomposition.
 In Section 2 we present the proposed online learning-decomposition framework.
-Section IV details the experimental setup, followed by Section V, where we assess the performance of the proposed framework.
+Section IV details the experimental setup, followed by Section V, where we assess the performance and model complexity of the proposed framework.
 Finally, Section VI presents our conclusions.
@@
 II-A SLA Decomposition and Risk Models
-An E2E SLA, denoted by $s_{e2e}$, is a collection of Service Level Objectives (SLOs) linked to specific performance indicators.
+An E2E SLA, denoted by $s_{e2e}$, is a collection of SLOs linked to specific performance indicators.
 The SLO vector outlines the SLA’s performance requirements in a sequential manner.
@@
 Particularly, the acceptance probability becomes time-dependent.
-At a discrete time step $t$, the acceptance probability (in a single domain) of the SLA request $s$ is denoted as $P_{t}(s)$, which is influenced by the state $\omega_{t}$ of the domain, as stated in Section II-B.
+At a discrete time step $t$, the acceptance probability (in a single domain) of the SLO request $s$ is denoted as $P_{t}(s)$, which is influenced by the state $\omega_{t}$ of the domain, as stated in Section II-B.
 Given that $\omega_{t}$ evolves over time, the acceptance probability $P_{t}(s)$ also varies, even for the same SLOs.
@@
 Figure 2: Illustration of RADE framework.
-In this section, we present the proposed framework Real-time Adaptative DEcomposition (RADE).
+In this section, we present the proposed framework Real-time Adaptive DEcomposition (RADE).
 III-A Two-step Decomposition with NN-based Risk Models
@@
 Following the design in [7], the base model is an NN.
-To account for monotonicity, we specifically choose Absolute Weight Transformation (AWET) model that shows prominent performance while requiring the least training effort, as suggested in [4].
+To account for monotonicity described in (4), we employ Absolute Weight Transformation (AWET) approach that shows prominent performance.
+AWET ensures that the weights remain non-negative, a sufficient condition for an NN to be monotonic, while still allowing the model’s weights to be optimized freely during training.
 Online update.
@@
 Algorithm 1 RADE framework
-By updating the risk models online, we ensure that the decomposition remains responsive and resilient to changes.
-This continuous adaptation process not only enhances the accuracy of the model but also improves the overall performance and reliability of the system.
 Alg.
 1 details the steps of the online update and decomposition mechanism.
-In the subsequent sections, we will present the simulation environment for evaluations, followed by empirical results demonstrating its effectiveness.
+The RADE framework operates in a time-stepped manner, where two key processes—learning and inference—run in parallel during each time step.
+At the beginning of each step, the framework receives feedback from the network regarding the performance of recent service requests.
+This feedback, represented as a set $\mathcal{K}_{t}$, is pushed into a FIFO memory buffer $\mathcal{M}$ to ensure that only the most recent and relevant data are stored.
+This buffer helps balance between old and new feedback.
+Using this stored feedback, the framework updates its risk models $\mathcal{F}_{\theta}$ through an online gradient descent process, adjusting the parameters based on the gradient of the loss function defined in (3).
+These continuously updated models help the system stay adaptive to changing network conditions.
+Simultaneously, the framework processes a set of new service requests $\mathcal{R}_{t}$, each accompanied by an E2E SLA $\mathcal{S}_{e2e}$ that defines the expected performance.
+The E2E SLAs are then decomposed into partial SLAs $\mathcal{S}_{partial}$ for individual network domains by leveraging the latest risk models and optimizing (1) and (2).
+The decomposition process ensures that the E2E SLA are allocated across the network domains in an optimal manner such that the overall E2E acceptance probability is maximized.
+In the subsequent sections, we will present the simulation environment for evaluations, followed by empirical results.
 IV Experimental Setup
@@
 V Results
+V-A Performance
 Fig.
@@
 RADE* shows improved performance over the Static method, which indicates that dynamically updating the risk models, even without the FIFO memory buffer, leads to a better adaptability than the Static approach.
-The RADE method, which includes the FIFO memory buffer for maintaining historical feedback, outperforms both Static and RADE*, highlighting the importance of the use of the FIFO buffer in enhancing the robustness and stability of the framework, particularly under varying traffic conditions.
+The RADE method, which includes the FIFO memory buffer for maintaining historical feedback, outperforms both Static and RADE*, highlighting the importance of the use of the FIFO buffer in enhancing the robustness and stability of the framework, particularly under low arrival rate conditions.
 It is interesting to see that the performance of RADE* and RADE becomes nearly identical at higher arrival rates because, under heavy traffic, the system receives a large volume of recent feedback.
@@
 Figure 5: Average E2E acceptance probability vs. corruption rate at arrival rate=0.5.
+V-B Complexity analysis
+The comparison of time complexity is provided in Table I.
+Three methods (Static, RADE* and RADE) are analyzed regarding both the complexity of base model training and the decomposition processes, which run in parallel.
+Assume the architecture shared by all methods has a training complexity $O(f(D))$ for a single iteration, and $O(g(D))$ for decomposition, where $D$ is the size of the input data, and $f$, $g$ are functions dependent on the architecture.
+For the Static method, the time complexity is simply $O(g(D))$, since there is no additional overhead incurred by updates.
+For RADE*, the total complexity for $N$-iteration training is $O(Nf(D))$, and the overall complexity is $O(\max(Nf(D),g(D)))$.
+Similarly, RADE behaves like RADE* but with the additional use of the FIFO memory buffer.
+Assume the FIFO memory buffer can hold up to $Q$ data points, the overall complexity is $O(\max(Nf(Q),g(D)))$.
+Typically, $Q$ is larger than $D$ to ensure effective utilization of the FIFO memory buffer.
+In summary, the Static method is the most efficient in terms of time complexity, followed by RADE*.
+RADE, however, is the most computationally intensive method, primarily due to the additional overhead of performing online training on data stored in the FIFO memory buffer.
+Method Time complexity
+Static $O(g(D))$
+RADE* $O(\max(Nf(D),g(D)))$
+RADE $O(\max(Nf(Q),g(D)))$
+TABLE I: Comparison of time complexity.
 VI conclusion
@@
 Empirical results demonstrate that our framework outperforms state-of-the-art static methods, offering a more resilient and accurate decomposition of SLAs, even with limited data availability.
-In future work, we will investigate the potential of Deep Reinforcement Learning (DRL) to further enhance the SLA decomposition process.
-DRL could be a promising choice for this task due to its ability to learn optimal long-term policies through interaction with dynamic environments, which makes it an attractive option for improving SLA management in next-generation networks.
+In future work, we will explore the potential of Deep Reinforcement Learning (DRL) to further refine the SLA decomposition process.
+DRL is particularly promising for this task due to its capacity to optimize long-term rewards, which is crucial in dynamic environments where immediate feedback may not fully capture the impact of decisions.
+By learning to cope with delayed rewards, DRL can develop strategies that anticipate the cumulative effects of actions over time, making it an ideal approach for managing SLA decomposition in complex, evolving network conditions.